OpenStudy (anonymous):

Find the value of c such that the point (2, 5) lies on the graph of f(x) = cx^2 + 4x − 2. c=????

OpenStudy (tkhunny):

Have you considered substituting the give values? x = 2, y = 5

OpenStudy (marigirl):

Yes write it as 5=c(2)^2+4(2)-2 and then solve for c.

OpenStudy (anonymous):

Would the answer be -5?

OpenStudy (anonymous):

-1/4?

OpenStudy (whpalmer4):

Let's find out! \[f(x) = -5x^2+4x-2\]\[f(2) = 5 = -5(2)^2+4(2)-2\]\[5 = -5(4) + 4(2) - 2\]\[5 = -20 +8 -2\]Is that true?

OpenStudy (whpalmer4):

Let's find out: \[5 = -\frac{1}{4}(2)^2+4(2) - 2\]\[5 = -1+8-2\]\[5 = 5\checkmark\]

OpenStudy (anonymous):

awesome! people are so helpful here

OpenStudy (whpalmer4):

You should go back and figure out what mistake you made that made you think \(c=-5\)... learn from your mistakes!